OpenStudy (anonymous):

||6x+2|-12|=7 , find the exact value for x?

OpenStudy (anonymous):

\[\left| \left| 6x+2 \right| -12\right|=7\] The rule you need is this: \[\left| f(x) \right|=5\] becomes \[f(x)=5\] and \[f(x)=-5\]

OpenStudy (anonymous):

i don't understand can you explain more

OpenStudy (anonymous):

Whenever an absolute value is undone the result is two equations that are equal to the positive and the negative of what the absolute was equal too.

OpenStudy (anonymous):

Therefore, \[\left| \left| 6x+2 \right| -12\right|=7\] Becomes \[\left| 6x+2 \right|-12=7\] AND \[\left| 6x+2 \right|=-7\]

OpenStudy (anonymous):

oops the final equation in the last post was wrong

OpenStudy (anonymous):

\[\left| 6x+2 \right|-12=7\] AND \[\left| 6x+2 \right|-12=-7\]

OpenStudy (anonymous):

Do you understand why we split the equation from one into two?

OpenStudy (anonymous):

yes and the how do i find x value?